The atmospheric pressure at the surface of a swimming pool is 101.3 kPa. What is the pressure 2.90 m below the surface? (In kPa, a kPa is a kiloPascal or 1000 N/m2)
Answer #1

Pressure at depth h below liquid surface is (hdg) more than pressure at surface. d is density of liquid. Hence pressure 2.9m below water surface

= P(surface) + hdg

= 101.3×10^3 Pa + (2.9m)(1000kg/m^3)(9.81m/s^2)

= 129.7×10^3 N/m^2 = 129.7 kPa